OpenStudy (anonymous):

Write a linear factorization of the function. f(x) = x4 + 4x2

OpenStudy (anonymous):

here are the answer choices.. f(x) = x2(2x + i)2 f(x) = x2(x + 2i)2 f(x) = x2 (x + 2i)(x - 2i) f(x) = x2 ( 2x + i)( 2x - i)